OpenStudy (loser66):

Is there a case of 2 distinct lines lie on the 2 different planes? Please, help

OpenStudy (loser66):

I meant if AB, CD are in plane1, is there the case of AB, CD are in plane2 which is different from plane1

imqwerty (imqwerty):

well if a line lies on x-y plane we wuld write it as- ax+by+c=0 nd if it is on x-z plane we wuld write it as ax+bz+c=0 so u can get the same eq....

OpenStudy (loser66):

I don't get it, please, explain more.

OpenStudy (loser66):

and what is nd ? on "nd if it is....."

imqwerty (imqwerty):

ok the standard form of a line present in x-y plane is- ax+by+c=0 u can write it as - -ax-by=c x/(c/-a)+y/(c/-b)=1 where c/-a = x intercept c/-b = y intercept

imqwerty (imqwerty):

similarly if u go with x-z axes

imqwerty (imqwerty):

u have the standard line equation as ax+bz+c=0 u can write it as- x/(c/-a) + z/(c/-b) =1 nd the intercepts culd be the same in both cases so the equations can be similar but they will differ cause they r jst in diff planes

OpenStudy (loser66):

So, it is false, right? there is not the case of if AB, CD are in plane 1, they cannot be in plane2 which is different from plane 1, right?

OpenStudy (loser66):

I appreciate your help, but it will be perfect if you use standard English instead of the "teen language" since I am not a native one. I cannot understand all of the teen words. :)